A constant force changes the speed of a 100kg sprinter from 4.0ms-1 in 1.0s calculate

A ) the acceleration of the sprinter during this time


B) the magnitude of the force
1
Expert's answer
2020-05-07T08:40:29-0400

(A) The acceleration of the sprinter

"a=\frac{\Delta v}{\Delta t}=\frac{4.0\:\rm m/s}{1.0 \:\rm s}=4.0\:\rm m/s^2."

(B) The magnitude of the force

"F=ma=(100\:\rm kg )\times (4.0 \:\rm m/s^2)=400\:\rm N."